# Combining Textual and Structural Information for Premise Selection in Lean

### v1

@@ Abstract
Premise selection is a key bottleneck for scaling theorem proving in large formal libraries. Yet existing language-based methods often treat premises in isolation, ignoring the web of dependencies that connects them. We present a graph-augmented approach that combines dense text embeddings of Lean formalizations with graph neural networks over a heterogeneous dependency graph capturing both state–premise and premise–premise relations. On the LeanDojo Benchmark, our method outperforms the ReProver language-based baseline by over $25\%$ across standard retrieval metrics. These results demonstrate the power of relational information for more effective premise selection.
@@ 1 Introduction and related work
Recent advances in artificial intelligence, particularly large language models (LLMs), have demonstrated increasing efficacy in formal mathematics and interactive theorem proving (Polu and Sutskever, 2020; Jiang et al., 2021; Xin et al., 2024). A central task in this domain is premise selection: retrieving relevant theorems and definitions from extensive libraries to guide the proofs of new theorems. Effective premise selection underpins many automated reasoning tools, such as Sledgehammer (Böhme and Nipkow, 2010) and modern AI-driven provers for Lean (Song et al., 2023).
LLM-based approaches such as ReProver (Yang et al., 2023) use dual-encoder models to map proof states and premises into a shared vector space, retrieving relevant premises via dense ByT5 embeddings. However, they ignore the structural relationships in formal libraries: the references and dependencies among axioms, theorems, lemmas, and definitions (entries), which offer strong prior knowledge for guiding proofs. Attempts to exploit this structure include (Bauer et al., 2023) and (Ferreira and Freitas, 2020), but the former neglects Lean 4’s state–tactic framework, while the latter tackles premise selection only in natural text.
@@ 2.1 Augmented data extraction from the Mathlib Library
Base dataset: proof states and premises In the LeanDojo machine learning framework (Yang et al., 2023), premise selection is formulated as the following task. For each proof state $s$, given its textual representation $x_{s}$, we aim to identify the list of relevant premises $y_{s}$ from the (Mathlib (The mathlib Community, 2020)) library that will be used in the next tactic. For example, for the proof state in Figure 3, the tactic uses the premise not_or. Although only one premise appears in this example, in general, a tactic application may use multiple premises or none. Let $S$ denote the set of proof states, and $X_{S}$ the text representations for each $s\in S$.
Let $P$ be the set of all potential premises available in the library. Each premise $p\in P$ has a library definition in Lean code, and we denote by $X_{P}$ the code (text) representations of the premises in $P$. For example, in Figure 1, the entire block defines the premise premise_example. The dataset in (Yang et al., 2023) can therefore be summarized as the tuple $(X_{P},X_{S},Y_{S})$, where $Y_{S}$ denotes the set of all lists of relevant premises $y_{s}$.
@@
By merging $V=P\cup S$ as the node set, $X=X_{S}\cup X_{P}$ as the textual representations of the corresponding Lean formalizations, and references $E\subseteq V\times\mathcal{R}\times V$ as the edge set, we can summarize a Lean library as a text-attributed directed graph $G=(V,X,E)$. Here $\mathcal{R}$ is the set of relation types described in the next paragraph. We extend LeanDojo’s open-source extraction framework (Yang et al., 2023) to query the proof assistant for this additional information and extract this full dataset, i.e., the tuple $(G,Y_{S})$. The dataset (for Mathlib) thus now includes $(X_{P},X_{S},Y_{S})$ from (Yang et al., 2023), but adds the additional premise-premise and premise-state edges, forming a directed graph.
@@
File and import graph To remain consistent with the framework of (Yang et al., 2023), we also use the separate directed graph of imports between the files and a map between entries, states, and the files where they were defined. We do not use this information during training, and it is not employed by the model. We leave the utilization of this information for future work. However, during evaluation, this graph is used to restrict the premise selection only to premises that would have been available to the model at that point in the file: premises from imports, and premises defined before the current entry. For complete details, see (Yang et al., 2023).
@@ 2.2 Dataset statistics
Table 2 in the appendix summarizes the graph statistics for the augmented LeanDojo benchmark extracted from the Mathlib library commit 29dcec074de168ac2bf835a77ef68bbe069194c5, the one used in the official repository of (Yang et al., 2023). This allows us to directly compare our results with the results of their approach.
@@ 3.1 GNN-augmented premise retrieval
Step 1: initial text embeddings. The textual representations of premises $x_{p}$ and states $x_{s}$ are initially embedded using ReProver’s ByT5 dual-encoder (Yang et al., 2023), yielding initial node feature vectors $\mathbf{h}_{p}^{(0)}$ and $\mathbf{h}_{s}^{(0)}$, respectively.
Step 2: GNN-based refinement. We employ a Relational Graph Convolutional Network (RGCN) (Schlichtkrull et al., 2018) to propagate information over the heterogeneous directed graph. Each premise node’s embedding $\mathbf{h}_{p}$ is updated iteratively using
@@ 3.2 Training objective
We use the InfoNCE (van den Oord et al., 2018; Rusak et al., 2025) loss, which for a given list $y_{s}$ of valid premises, calculates as:
@@
where the hyperparameter $\tau$ is a scalar temperature. Finally, $f(s,p)=\mathrm{sim}(\mathbf{h}_{s}^{(L)^{\prime}},\mathbf{h}_{p}^{(L)})$. The InfoNCE loss contrasts positive and negative premises for each state, summing probabilities over multiple valid premises. The GNN is trained transductively on the full premise graph, excluding all edges from proof relations (the third item in the enumeration of relation types in Section 2.1) to prevent trivial memorization and promote learning of general structural patterns.
@@ 3.3 Other optimizations
We use an ensemble of six independently-trained models (N=6), averaging their outputs to form the final prediction. This approach mitigates initialization sensitivity, which is a known issue in GNNs (Li et al., 2023). In addition, we apply exponential model averaging (EMA) to further improve performance and generalization (Morales-Brotons et al., 2024).
@@ 4.1 Experimental setup
For our experiments, we use the LeanDojo Mathlib benchmark dataset with graph augmentation. We adopt the same train/validation/test split as (Yang et al., 2023) (the “random” split) and evaluate with the same metrics: Recall@1, Recall@10, and Mean Reciprocal Rank (MRR). Our baseline is the ReProver retriever with a ByT5-small encoder. We tune our model’s hyperparameters using Optuna on the validation set (see Appendix B). The number of GNN layers used is $L=2$.
@@ 4.2 Results
Our GNN-augmented retriever outperforms the baseline across all metrics. Note that the baseline results differ from those reported in (Yang et al., 2023). On GitHub [^1] https://github.com/lean-dojo/ReProver/discussions/51, the authors explain that this is potentially due to improvements made to the ReProver system from the time of their publication.
@@ 5 Conclusion
We introduced a graph-augmented language approach to premise selection in Lean. By extracting fine-grained syntactic dependencies and propagating structural information via a GNN, our method produces embeddings that outperform text-based baselines. Future work includes exploring advanced GNN architectures, such as graph attention networks, to better incorporate structural information. We can also analyze the effect of the number of GNN layers. Finally, the model will be evaluated on more realistic splits, such as the LeanDojo "novel" split or a split based on the creation time of premises.
@@ 6 Acknowledgments
The authors acknowledge the financial support of the Slovenian Research Agency via the Gravity project AI for Science, GC-0001, and of the Renaissance Philanthropy via the project Bridging AI, Proof Assistants, and Mathematical Data (BRIDGE). We are especially grateful to Boshko Koloski for the beneficial discussion related to the training of our model.
@@ B.1 Hyperparameter tuning
The hyperparameters for our GNN retrieval model were determined through a two-stage search using the Optuna framework [Akiba et al., 2019]. The first stage focused on architectural choices, such as GNN layer type, hidden dimensions, learning rate, and model structure (e.g., separate vs. shared GNNs for premises and contexts). The objective was to maximize the Recall@10 metric on the training set for the first batch.

### v2

@@ Abstract
Premise selection is a key bottleneck for scaling theorem proving in large formal libraries. Yet existing language-based methods often treat premises in isolation, ignoring the web of dependencies that connects them. We present a graph-augmented approach that combines dense text embeddings of Lean formalizations with graph neural networks over a heterogeneous dependency graph capturing both state–premise and premise–premise relations. On the LeanDojo Benchmark, our method outperforms the ReProver language-based baseline by over $25\%$ across standard retrieval metrics. These results suggest that relational information is beneficial for premise selection.
@@ 1 Introduction and related work
Recent advances in artificial intelligence, particularly large language models (LLMs), have demonstrated increasing efficacy in formal mathematics and interactive theorem proving (polu2020generative; jiang2021lisa; xin2024advancing). A central task in this domain is premise selection: retrieving relevant theorems and definitions from extensive libraries to guide the proofs of new theorems. Effective premise selection underpins many automated reasoning tools, such as Sledgehammer (bohme2010sledgehammer) and modern AI-driven provers for Lean (song2023towards).
LLM-based approaches such as ReProver (yang2023leandojo) use dual-encoder models to map proof states and premises into a shared vector space, retrieving relevant premises via dense ByT5 embeddings. However, they ignore the structural relationships in formal libraries: the references and dependencies among axioms, theorems, lemmas, and definitions (entries), which offer strong prior knowledge for guiding proofs. Attempts to exploit this structure include (10.5555/3666122.3668329) and (ferreira-freitas-2020-premise), but the former neglects Lean 4’s state–tactic framework, while the latter tackles premise selection only in natural text.
@@ 2.1 Augmented data extraction from the Mathlib Library
Base dataset: proof states and premises In the LeanDojo machine learning framework (yang2023leandojo), premise selection is formulated as the following task. For each proof state $s$, given its textual representation $x_{s}$, we aim to identify the list of relevant premises $y_{s}$ from the (Mathlib (mathlib)) library that will be used in the next tactic. For example, for the proof state in Figure 3, the tactic uses the premise not_or. Although only one premise appears in this example, in general, a tactic application may use multiple premises or none. Let $S$ denote the set of proof states, and $X_{S}$ the text representations for each $s\in S$.
Let $P$ be the set of all potential premises available in the library. Each premise $p\in P$ has a library definition in Lean code, and we denote by $X_{P}$ the code (text) representations of the premises in $P$. For example, in Figure 1, the entire block defines the premise premise_example. The dataset in (yang2023leandojo) can therefore be summarized as the tuple $(X_{P},X_{S},Y_{S})$, where $Y_{S}$ denotes the set of all lists of relevant premises $y_{s}$.
@@
By merging $V=P\cup S$ as the node set, $X=X_{S}\cup X_{P}$ as the textual representations of the corresponding Lean formalizations, and references $E\subseteq V\times\mathcal{R}\times V$ as the edge set, we can summarize a Lean library as a text-attributed directed graph $G=(V,X,E)$. Here $\mathcal{R}$ is the set of relation types described in the next paragraph. We extend LeanDojo’s open-source extraction framework (yang2023leandojo) to query the proof assistant for this additional information and extract this full dataset, i.e., the tuple $(G,Y_{S})$. The dataset (for Mathlib) thus now includes $(X_{P},X_{S},Y_{S})$ from (yang2023leandojo), but adds the additional premise-premise and premise-state edges, forming a directed graph. The modified extraction code and the learning pipeline used is available at https://github.com/JobPetrovcic/GNNReProver/tree/lighweight.
@@
File and import graph To remain consistent with the framework of (yang2023leandojo), we also use the separate directed graph of imports between the files and a map between entries, states, and the files where they were defined. We do not use this information during training, and it is not employed by the model. We leave the utilization of this information for future work. However, during evaluation, this graph is used to restrict the premise selection only to premises that would have been available to the model at that point in the file: premises from imports, and premises defined before the current entry. For complete details, see (yang2023leandojo).
@@ 2.2 Dataset statistics
Table 3 in the appendix summarizes the graph statistics for the augmented LeanDojo benchmark extracted from the Mathlib library commit 29dcec074de168ac2bf835a77ef68bbe069194c5, the one used in the official repository of (yang2023leandojo). This allows us to directly compare our results with the results of their approach.
@@ 3.1 GNN-augmented premise retrieval
Step 1: initial text embeddings. The textual representations of premises $x_{p}$ and states $x_{s}$ are initially embedded using ReProver’s ByT5 dual-encoder (yang2023leandojo), yielding initial node feature vectors $\mathbf{h}_{p}^{(0)}$ and $\mathbf{h}_{s}^{(0)}$, respectively.
Step 2: GNN-based refinement. We employ a Relational Graph Convolutional Network (RGCN) (schlichtkrull2018rgcn) to propagate information over the heterogeneous directed graph. Each premise node’s embedding $\mathbf{h}_{p}$ is updated iteratively using
@@ 3.2 Training objective
We use the InfoNCE (infonce; rusak2024infonce) loss, which for a given list $y_{s}$ of valid premises, calculates as:
@@
where the hyperparameter $\tau$ is a scalar temperature. Note that by minimizing this loss, we train a model of the scoring function $f(s,p)=\mathrm{sim}(\mathbf{h}_{s}^{(L)^{\prime}},\mathbf{h}_{p}^{(L)})$. The InfoNCE loss contrasts positive and negative premises for each state, summing probabilities over multiple valid premises. The GNN is trained transductively on the full premise graph, excluding all edges from proof relations (the third item in the enumeration of relation types in Section 2.1) to prevent trivial memorization and promote learning of general structural patterns.
@@ 3.3 Other optimizations
We use an ensemble of six independently-trained models (N=6), averaging their outputs to form the final prediction. This approach mitigates initialization sensitivity, which is a known issue in GNNs (gnn_init). In addition, we apply exponential model averaging (EMA) to further improve performance and generalization (ema). Finally, compared to LLMs, GNNs use relatively less memory, removing the need to sample negative examples. We thus simply use all other premises in the library as negatives.
@@ 4.1 Experimental setup
For our experiments, we use the LeanDojo Mathlib benchmark dataset with graph augmentation. We adopt the same train/validation/test split as (yang2023leandojo) (the “random” split) and evaluate with the same metrics: Recall@1, Recall@10, and Mean Reciprocal Rank (MRR). Our baseline is the ReProver retriever with a ByT5-small encoder. We tune our model’s hyperparameters using Optuna on the validation set (see Appendix B). The number of GNN layers used is $L=2$.
@@ 4.2 Results
Results reported in Table 1 show that our GNN-augmented retriever outperforms the baseline across all metrics. Note that the baseline results differ from those reported in (yang2023leandojo). On GitHub [^1] https://github.com/lean-dojo/ReProver/discussions/51, the authors explain that this is potentially due to improvements made to the ReProver system from the time of their publication.
@@
4.3 Ablations
To understand the contribution of different components in our model, we perform an ablation study on the augmented LeanDojo data set. Specifically, we examine the effect of excluding the graph between contexts and premises and the graph between premises and premises. Table 2 summarizes the results of the ablations. Note that due to computational cost, no ensembling was used in all cases for fair comparison and thus the results for No ablation in Table 2 do not match those in Table 1.
Model R@1 R@10 MRR
No ablation 17.43% 48.52% 0.4010
No ablation + EMA 17.74% 48.70% 0.4048
Context graph ablation 17.30% 49.99% 0.4008
Context graph ablation + EMA 17.58% 50.13% 0.4057
Premise graph ablation 17.45% 49.30% 0.4008
Premise graph ablation + EMA 18.18% 49.98% 0.4096
Table 2: Ablation study on the LeanDojo test set.
After performing ablations, we realized that removing parts of the dependency graph even improves performance. This leads to the possibility that the utilization of the graph structure is not the main factor behind the superior performance, but rather that this is caused by the different choices of the loss function and sampling strategy. We leave this scrutinization for future work.
@@ 5 Conclusion
We introduced a graph-augmented language approach to premise selection in Lean. By extracting fine-grained syntactic dependencies and propagating structural information via a GNN, our method produces embeddings that outperform text-based baselines.
As the ablation section suggests, however, the superior performance might be attributed to the different training paradigm choices rather than the model’s utilization of graph structure. Besides further analysis of the root cause of the improvement, future work includes exploring advanced GNN architectures, such as graph attention networks, to better incorporate structural information. Finally, the model will be evaluated on more realistic splits, such as the LeanDojo "novel" split or a split based on the creation time of premises.
@@ 6 Acknowledgments
The authors acknowledge the financial support of the Slovenian Research Agency via the Gravity project AI for Science, GC-0001 and via the research core funding No. P2-0103, the Air Force Office of Scientific Research under award number FA9550-21-1-0024, and the Renaissance Philanthropy via the project Bridging AI, Proof Assistants, and Mathematical Data (BRIDGE). We are especially grateful to Boshko Koloski for the beneficial discussion related to the training of our model.
@@ B.1 Hyperparameter tuning
The hyperparameters for our GNN retrieval model were determined through a two-stage search using the Optuna framework [optuna]. The first stage focused on architectural choices, such as GNN layer type, hidden dimensions, learning rate, and model structure (e.g., separate vs. shared GNNs for premises and contexts). The objective was to maximize the Recall@10 metric on the training set for the first batch.